Which of the ordered pairs is a solution for the equation 4x – 2y = 8?
A) (0, 4) B) (–2, 0) C) (–2, –4) D) (0, –4)

Answer by Edwin McCravy(20086)

Which of the ordered pairs is a solution for the equation 4x – 2y = 8?


A) (0, 4)

Let's substitute (0) for x, and (4) for y in 
4x - 2y = 8 and see if we get the same number 
on both sides of the equal sign:

    4x - 2y = 8

4(0) - 2(4) = 8
      0 - 8 = 8
         -8 = 8

No, -8 and 8 are not the same number.



B) (–2, 0) 

Let's substitute (-2) for x, and (0) for y in 
4x - 2y = 8 and see if we get the same number 
on both sides of the equal sign:

     4x - 2y = 8

4(-2) - 2(0) = 8
      -8 - 0 = 8
          -8 = 8

No, -8 and 8 are not the same number.


C) (–2, –4)

Let's substitute (-2) for x, and (-4) for y in
4x - 2y = 8 and see if we get the same number 
on both sides of the equal sign:

    4x - 2y = 8

4(-2) - 2(-4) = 8
       -8 + 8 = 8
            0 = 8

No, 0 and 8 are not the same number.


 D) (0, –4)

Let's substitute (0) for x, and (-4) for y in
4x - 2y = 8 and see if we get the same number 
on both sides of the equal sign:

     4x - 2y = 8

4(0) - 2(-4) = 8
       0 + 8 = 8
           8 = 8

Yes, 8 and 8 are the same number! So the 
correct answer is the last one, D).

Edwin